A 3kg mass with an initial velocity of 5m/s collides with and stick to a 2kg mass with an initial velocity of-3m/s .find the final velocity of the composite mass
1
Expert's answer
2020-07-06T15:13:49-0400

The total momentum in any closed system will remain constant. So, we have


"m_1v_1+m_2v_2=(m_1+m_2)v \to v=\frac{m_1v_1+m_2v_2}{m_1+m_2}=\frac{3\cdot 5+2\cdot (-3)}{3+2}=1.8m/s"